OpenStudy (handsomechickenlover):

What is the slope of the line whose equation is −48=2x−8y? Medal will be given!

jabez177 (jabez177):

Alright let's just turn the equation around.

OpenStudy (handsomechickenlover):

2x-8y=-48?

jabez177 (jabez177):

Wait... Let me see LOL Haven't quite mastered this yet...

OpenStudy (sooobored):

solve for y, the slope will be "m" \[y=mx+b\]

jabez177 (jabez177):

^ That's what I'm trying to do. So don't you turn the equation around?

OpenStudy (sooobored):

doesnt really matter

OpenStudy (sooobored):

i mean, you could write it b+mx=y slope will still be m

jabez177 (jabez177):

So how do we find it in this equation?

OpenStudy (studygurl14):

-48 = 2x - 8y Rearrange to y = mx +b -48 + -2x = -8y Divide both sides by -8

OpenStudy (handsomechickenlover):

-40+-2x=y?

OpenStudy (sooobored):

-48/-8 = 6 -2/-8 = 1/4 -- yo ucan just leave this as a fraction

OpenStudy (handsomechickenlover):

oh

OpenStudy (studygurl14):

Yep so you get 6 + (1/4)x = y So in y = mx + b, form y = (1/4)x + 6 m = slope

OpenStudy (handsomechickenlover):

Okay I get now thank you all :D

jabez177 (jabez177):

So do I! Thanks ahah!!